### v1

# Large Language Model

Abstract
The event-based Vision-Language Model (VLM) recently has made good progress for practical vision tasks. However, most of these works just utilize CLIP for focusing on traditional perception tasks, which obstruct model understanding explicitly the sufficient semantics and context from event streams. To address the deficiency, we propose EventVL, the first generative event-based MLLM (Multimodal Large Language Model) framework for explicit semantic understanding. Specifically, to bridge the data gap for connecting different modalities semantics, we first annotate a large event-image/video-text dataset, containing almost 1.4 million high-quality pairs of data, which enables effective learning across various scenes, e.g., drive scene or human motion. After that, we design Event Spatiotemporal Representation to fully explore the comprehensive information by diversely aggregating and segmenting the event stream. To further promote a compact semantic space, Dynamic Semantic Alignment is introduced to improve and complete sparse semantic spaces of events. Extensive experiments show that our EventVL can significantly surpass existing MLLM baselines in event captioning and scene description generation tasks. We hope our research could contribute to the development of the event vision community.
1 Introduction
Event cameras are sensors that asynchronously measure changes in intensity at each pixel with microsecond-level precision. Compared to traditional cameras, event cameras offer several significant advantages [14], including a high dynamic range ($>120$ dB), resistance to motion blur, high temporal resolution, and extremely low power consumption.These advantages have led to significant successes of event-based methods in different domains, such as robotics [18] and imaging applications [44, 36]. Even though event cameras offer richer semantic information—including continuous temporal motion and higher dynamic range lighting, exploration into high-level and open-world understanding tasks remains limited.
As VLMs [26, 35] achieve great successes in image understanding, the event-based community began to develop event-based VLMs for a comprehensive understanding of event data, which can be beneficial to downstream tasks [52, 48] like open-vocabulary segments [25] for large-scale scene understanding. Most of these works focus on image-event-text pair alignment based on CLIP [34] by bridging the modality gap between these data according to task scenarios. Though successful, these CLIP-based methods limit themselves in traditional perception task [51, 43] due to a coarse understanding of event data, which makes it hard to generalize to a fine-grained dynamic semantic task such as the detailed description for object motion pattern and the various relation between objects. It leads to the bottleneck in spatial understanding of many scenarios, for instance, autonomous driving or navigation. Recently, current multi-modal large language models (MLLMs) [29, 1] have utilized images or videos as inputs to obtain an accurate and fine-grained text description, lifting visual perception to natural-language-based comprehension of the world. From our perspective, MLLMs have the potential to overcome the scene understanding limitations of event data and offer a versatile language interface for user interaction.
However, to train an event-based MLLM, we are faced with two main challenges. First, a lack of high-quality text annotation in current image-event-text pair datasets hinders the model from capturing fine-grained semantic information [23, 40]. In those datasets, the coarse text descriptions like “This is a {category}” limit the model to the learning of category shapes and hinder the model from excavating more diverse event-based category knowledge such as colour or materials, which prohibits the fine-grained semantic understanding of event data. Second, different from images, event data, as a format of spatiotemporal point clouds, is incompatible with the current RGB encoder architecture. Moreover, preserving the highly spatiotemporal correlations within event data during feature extraction is challenging. An inadequate representation of events can impair event-image-text fine-grained alignment, resulting in sub-optimal scene understanding.
To solve these challenges mentioned above, we propose EventVL, the first event-based MLLM framework, as shown in Figure 2. As for data engineering, we first annotate almost 1.4 million high-quality image-text-event paired data by utilizing current powerful open-source MLLM models [7], which has surpassed the commercial MLLMs like GPT-4o  [1] in many understanding benchmarks. As for EventVL, we propose Event Spatiotemporal Representation to further excavate the spatiotemporal correlation within event data. We spatially and temporally segment event data into different frames and leverage attention blocks to capture the semantic interaction, which can enhance feature expression. To fine-grain align the latent feature space for images and events, we propose the Dynamic Semantic Alignment for rigorous high-resolution semantics extraction, enabling fine-grained alignment between events and images.
Figure 1: The Designed Description Generation Engine. For different kinds of datasets, we pre-define the corresponding problem list and prompts, which “Static Images” is for N-ImageNet [23], N-Caltech101 [31], “Human Motions” is for HARDVS [40] and “Drive Scenes” is for DSEC [16]. The coarse captions generated by MLLMs also are checked manually by random sampling. The pairs data marked as bad descriptions will be used to regenerate descriptions again with modified problems.
Comprehensive experiments verify that our proposed EventVL surpasses other SOTAs in zero/few-shot event captioning, event stream description generation tasks, and event-based dialogue. Furthermore, with a tiny number of parameters (almost 2.3B) compared to other MLLMs [30, 38], our model enables a low-cost for deployment to the real world. As shown in Figure LABEL:fig:abstract (a), EventVL can fully understand semantics and generate precise descriptions for event streams across diverse domains. Additionally, it supports multi-turn dialogues based on event data, enabling deeper semantic understanding, as shown in Figure LABEL:fig:abstract (b). In summary, our contribution has three folds:
(1) We present the First Event-based MLLM Framework, named EventVL, which aligns the large-scale event encoder with the image encoder and LLM. The model demonstrates strong performance on event-based generative tasks such as description or dialogue.
(2) We propose a simple but efficient description generation engine to get high-quality multi-modalities data by utilizing open-source VLMs. The processed datasets contain almost 1.4 million paired data across various domains. As far as we know, it is the biggest multi-modalities pair dataset in the event-based community.
(3) Event Spatiotemporal Representation is proposed for feature adaptive and efficient aggregation. We also propose a Dynamic Semantic Alignment module for fine-grained feature alignment, resulting in a precise and comprehensive event-based understanding.
2 Related Work
Event-based VLMs: Unlike the well-established field of image understanding, the event community is in the early stages of developing universal models for event comprehension. Initial efforts [24, 47] focused on using self-supervision to create a pre-trained backbone for event understanding. Inspired by CLIP’s success in computer vision, researchers began to explore its application for detailed and fine-grained event analysis. [43] suggested transforming event streams into an RGB-style representation to achieve event-text alignment using CLIP’s architecture. Recently, CLIP-based approaches have been applied to more complex downstream tasks such as motion classification [40] and scene segmentation [25]. However, these studies have mainly concentrated on discriminative tasks, falling short of achieving fine-grained understanding, which restricts their application in generative tasks like interactive event-conditioned text generation. In response to the increasing need for a nuanced understanding of event streams, we are committed to introducing a unified event-based MLLM framework. This framework is designed to fully comprehend the semantics of event streams through explicit language and is compatible with various types of event data.
Open-Source MLLMs: The emergence of open-source MLLMs [29, 17, 38, 30] has profoundly impacted the field of Artificial General Intelligence (AGI) by bolstering the ability to process both visual and textual data. In the past year, several open-source MLLMs have gained prominence such as MiniGPT-4 [53] and VisionLLM [39]. However, these models typically rely on traditional images or videos as input, which can result in the generation of low-quality images under extreme imaging conditions, thereby impairing comprehensive world understanding. Event cameras, with their distinct properties, can effectively address these challenges. Therefore, we are exploring the integration of event-based data into existing MLLMs to unlock the full semantic potential of event data.
3 Data Engineering
Figure 2: Overview of our proposed EventVL. EventVL mainly consists of two parts: (a) Event Spatiotemporal Representation and (b) Dynamic Semantic Alignment. First, we introduce Event Spatiotemporal Representation for extracting diverse event information comprehensively, which consists of two split techniques –(i) Hierarchical Temporal Split and (ii) Adaptive Spatial Split. Then, we utilize Dynamic Semantic Alignment to further push the image and event feature space in fine-grained alignment.
A high-quality dataset is crucial for training a MLLM. Some datasets pairing images or videos with events [23, 40] offer only basic categorical text annotations, akin to ”This is a category”, which lack detail and variety. While certain studies suggest using coarse image-text data to address the limited diversity of image-event datasets like N-ImageNet, these annotations often have a weak connection to the images and include much extraneous information. Our experiments indicate that training with such datasets leads to a decline in performance. Therefore, it is imperative to re-annotate this data to create a more precise and granular dataset, thereby enhancing the MLLM’s comprehension.
As shown in Figure 1, we design a simple but effective generation engine for different domain datasets that contain static images, dynamic scenes and human motions. We chose the InternVL2-76B [7] as the foundation model for description generation, which has shown its superiority and surpasses several SOTAs such as GPT-4o or Claude 3.5 on various understanding benchmarks. Utilizing it to label our data can reduce multiple costs compared to commercial engines and bring higher-quality descriptions.
Coarse Generation. We manually design the prompts and problem lists corresponding to various datasets, which can be referred to in our appendix. For image-event datasets (N-ImageNet, N-Caltech101), we first randomly sample questions from the problem list, then directly send images with questions into the InternVL2-76B for generating coarse captions. For video-event datasets (HARDVS, DSEC), it is inefficiency that utilizes all frames for generating description, which causes a large amount of computational consumption [25, 8]. To effectively extract long-term information from origin videos, we sampled uniformly $n$ frames. If the video contains less than $n$ frames, we adopt all frames. For balancing the cost and performance, we chose $n=14$. Then, the video and the correspondiong instruct is sent to InternVL2-76B for getting detailed description.
Manual Check. The cost of manually checking every image/video-text description is enormous. Hence, we develop the sampling strategies based on the category classification of each dataset for checking. For N-ImageNet, N-Caltech101, HARDVS and DSEC, they contain 1000 classes, 101 classes, 300 classes and 20 scenes respectively. We sample five coarse descriptions from each class or scene for checking. If the output description is not so good, all images/videos from this corresponding class or scenes will be sent to MLLM to regenerate the captions again with modified prompts and designated problems. Finally, all annotated multimodal pair datasets except N-Caltech101 will be mixed for model training. The dataset performance is validated as shown in Table 4. Note that our prepared dataset has almost 1.4 million image/video-event-text pairs covering diverse domains, which guarantees the event understanding performance.
4 Method
The overall workflow of EventVL is as shown in Figure 2, which is composed of a frozen image encoder, a trainable event encoder, and a frozen text decoder. Given batch-wise image-event-text pair data $\{(x_{im},x_{ev},\{x^{l}_{t}\}_{l=1}^{L})\}$, where event streams $x_{ev}$ contains $N_{\epsilon}$ event points, $x_{im}$ and $\{x^{l}_{t}\}_{l=1}^{L}$ denotes the image and text token list. We introduce Event Spatiotemporal Representation containing fine-grained spatiotemporal event semantic information, which is obtained by two split techniques – Hierarchical Temporal Split and Adaptive Spatial Split. After that, Dynamic Semantic Alignment is proposed for further projecting event data on image feature space in a compact latent space. We use the red-blue color map as the event representation consistent with recent approaches [46, 25]. This choice minimizes the gap between the event representation and the natural images for pre-trained models, thereby simplifying cross-modal alignment.
4.1 Event Spatiotemporal Representation
Event streams carry rich information about the captured scene [37]. To enable effective scene understanding, it is crucial to develop a comprehensive representation of event streams that aligns with neural networks for encoding spatiotemporal semantics. Hence, we propose Event Spatiotemporal Representation which divides and captures fine-grained temporal and spatial information. Hierarchical Temporal Split and Adaptive Spatial Split are proposed to obtain different groups of event frames containing temporal and spatial information, respectively. Finally, all event frames are concatenated, gathering the fine-grained spatiotemporal information.
Hierarchical Temporal Split. Due to the temporal dynamics in event data, there is a critical need for a novel approach to extract rich temporal semantics. Our method diverges from prior studies that have opted for introducing a multitude of parameters [52]. Instead, we propose Hierarchical Temporal Split to aggregate event temporal information in various levels, without the need for any additional parameters. As shown in Figure 2 (i), given an event stream $x_{ev}$ containing $N_{\epsilon}$ event points, we firstly define the event points per frame $\lambda n_{\epsilon},\lambda\in\{1,2\}$ for splitting the overall event stream into several groups, and transform these groups into RGB-style representations. After splitting hierarchically, we expand the events as multi-level events $\hat{\bm{x}}_{ev}=\{\{\hat{x}^{r}_{ev1}\}^{N_{1}}_{r=1},\{\hat{x}^{r}_{ev2}\}^{N_{2}}_{r=1},\hat{x}_{ev3}\}$ with three levels contains $N_{1}$, $N_{2}$, and $1$ RGB frames, respectively, where $N_{1}=\frac{N_{\epsilon}}{n_{\epsilon}}$, $N_{2}=\frac{N_{\epsilon}}{2n_{\epsilon}}$, and $\hat{x}^{r}_{ev*}\in\mathbb{R}^{H\times W}$ denotes the transformed RGB-style representations with a height of $H$ and a width of $W$. The level-3 event representation $\hat{x}_{ev3}$ aggregates all event points. In implementation, we set the $N_{\epsilon}$ as the fixed number and use padding for event streams containing less than $N_{\epsilon}$ event points. Finally, these temporal splits are sent to Adaptive Spatial Split for further spatial semantics extraction.
Adaptive Spatial Split. Previous research has focused on encoding complete event representations and aligning them with corresponding images within the feature space. While this method has led to improved model performance [48, 46], it requires resizing event frames to a low resolution, such as $224\times 224$. This resizing results in a significant loss of spatial information, particularly for high-resolution events like those in N-ImageNet [23], which can measure $480\times 640$. This issue worsens under poor imaging conditions, such as overexposure or rolling shutter effects, leading to even greater information loss. Hence, Adaptive Spatial Split is introduced for exploring full spatial event-based category information, as shown in Figure 2 (ii), which adaptively splits event representation or images into multiple patches uniformly in a high-resolution manner.
To preserve natural aspect ratios during processing, we adaptively select the most suitable aspect ratio from a predefined set $\mathcal{K}$, which encompasses all possible aspect ratio combinations derived from $n_{min}$ to $n_{max}$ tiles. We set $n_{min}$ to 1 and set $n_{max}$ to 6. For each input image or event representation, we determine its aspect ratio and compare it with the predefined ratios by calculating the absolute difference. In cases where multiple predefined aspect ratios are close matches (e.g., 1:1 and 2:2), we prioritize the one that does not exceed twice the area of the event representation to avoid significant enlargement. Subsequently, we divide the event representation into patches according to this chosen ratio and resize these patches to the desired dimensions. In detailed, given the multi-level events $\hat{\bm{x}}_{ev}=\{\{\hat{x}^{r}_{ev1}\}^{N_{1}}_{r=1},\{\hat{x}^{r}_{ev2}\}^{N_{2}}_{r=1},\hat{x}_{ev3}\}$, we only apply Adaptive Spatial Split to the level-3 event frame $\hat{x}_{ev3}$ and obtain multiple event patches spatially $\{\{\hat{x}^{r}_{ev3}\}^{N_{p}}_{r=1}\}$ ($N_{p}=6$). We concatenate multiple event patches with the origin multi-level events, obtaining Event Spatiotemporal Representation $\tilde{\bm{x}}_{ev}=\{\{\hat{x}^{r}_{ev1}\}^{N_{1}}_{r=1},\{\hat{x}^{r}_{ev2}\}^{N_{2}}_{r=1},\hat{x}_{ev3},\{\hat{x}^{r}_{ev3}\}^{N_{p}}_{r=1}\}$.
4.2 Dynamic Semantic Alignment
In this section, we aim to align event and image within the same latent space. We begin by inputting the event representation $\tilde{\bm{x}}{ev}$ into the event encoder, which is composed of multiple attention blocks. This process yields event embeddings $\{\{\phi^{r}_{ev1}\}^{N_{1}}_{r=1},\{\phi^{r}_{ev2}\}^{N_{2}}_{r=1},\phi_{ev3},\{\phi^{r}_{ev3}\}^{N_{p}}_{r=1}\}$ that are enriched with exchanged spatiotemporal knowledge and enhanced semantics. For feature alignment, we select $\bm{\phi}_{ev}=\{\phi_{ev3},\{\phi^{r}_{ev3}\}^{N_{p}}_{r=1}\}$. To obtain the corresponding image embeddings for alignment, we can perform the Adaptive Spatial Split for image data $x_{im}$ and get the segmented image patches concatenated with $x_{im}$ to obtain $\{x_{im},\{x^{r}_{im}\}^{N_{p}}_{r=1}\}$, which is inputted into the image encoder to obtain image embeddings $\bm{\phi}_{im}=\{\phi_{im},\{\phi^{r}_{im}\}^{N_{p}}_{r=1}\}$. Following the feature extraction by the attention blocks, each frame in $\bm{\phi}{ev}$ is sufficiently endowed with global temporal semantics and enhanced spatial knowledge. Finally, we conduct a cosine alignment between $\bm{\phi}_{ev}$ and $\bm{\phi}_{im}$.
Compared to ordinary images, event representation lacks detailed surface descriptions, such as material and texture. It motivates us to project events into the same compact feature space as images, thereby implicitly learning missing attribute descriptions. Hence, we leverage simple cosine similarity loss to maintain their mutual information,
$\vspace{-0.1cm}\mathcal{L}_{c}=1-\mathop{\textnormal{cos}}(\bm{\phi}_{ev},\bm{\phi}_{im})=1-\frac{\bm{\phi}_{ev}\cdot\bm{\phi}_{im}}{\|\bm{\phi}_{ev}\|\|\bm{\phi}_{im}\|}.$ (1)
By applying this loss, spatiotemporal fine-grained alignment can be achieved, and event data are projected into the same latent space as image data.
Dataset Type Resolution Split Category Scale
N-ImageNet [23] Image $480\times 640$ train 800 1,023,907
val 200 257,260
HARDVS [40] Video $346\times 260$ train 240 61,349
val 60 13,915
DSEC [16] Video $640\times 480$ train 42 2,173
N-Caltech101 [31] Image $302\times 245$ train 80 29,550
val 21 5,290
Total - - - 1,443 1,393,444
Table 1: Statistical information about datasets annotated by Section 3. Each data pair contains an accurate description.
4.3 Framework Optimization
Generally, maximizing the likelihood function below to align the representation spaces of image/video and text is a widely-used approach for pre-training [29, 45]. We also leverage this pattern to our event-text alignment. For a given event representation embeddings $\bm{\phi}_{ev}$ and a conversation list of $L$ text tokens $\bm{x}_{t}=\{x^{1}_{t},x^{2}_{t},...,x^{L}_{t}\}$, the likelihood of this list can be written as follows:
$\vspace{-0.1cm}p(\bm{x}_{t}\mid\bm{\phi}_{ev},\bm{x}_{\textnormal{instruct}})=\prod_{l=1}^{L}\phi_{t}(x_{t}^{l}\mid\bm{\phi}_{ev},\bm{x}_{\textnormal{instruct}},x_{t}^{1:l-1}),$ (2)
where $\bm{x}_{\textnormal{instruct}}$ is the instruction token, and $\phi_{t}(\cdot)$ is the text decoder. Then we perform event-text alignment by minimizing the negative log-likelihood, as follows:
$\vspace{-0.1cm}\mathcal{L}_{ev,t}=-\mathop{\textnormal{log}}p(\bm{x}_{t}\mid\bm{\phi}_{ev},\bm{x}_{\textnormal{instruct}}).$ (3)
We also have observed that text inputs often include detailed descriptions such as texture and color, which are not present in the corresponding event data due to the nature of event imaging. Relying solely on event-text alignment can lead to suboptimal outcomes, as it may focus on missing semantic elements. Therefore, to refine the embedding alignment, we incorporate the image embedding as a prior in Equation 3, effectively limiting the search space and enhancing the overall alignment process:
$\vspace{-0.1cm}\mathcal{L}_{ev,im,t}=-\mathop{\textnormal{log}}p(\bm{x}_{t}\mid\frac{1}{2}(\bm{\phi}_{ev}+\bm{\phi}_{im}),\bm{x}_{\textnormal{instruct}}).\vspace{-0.1cm}$ (4)
The overall training objectives can be written as follows:
$\vspace{-0.1cm}\mathcal{L}=\frac{1}{2}\lambda_{1}(\mathcal{L}_{ev,t}+\mathcal{L}_{ev,im,t})+\lambda_{2}\mathcal{L}_{c}.\vspace{-0.1cm}$ (5)
Note that the loss is computed in a batch and we omit the batch notation for clarity.
Dataset Model #param Bleu-1 Bleu-2 Bleu-3 Bleu-4 METEOR ROUGE-L CIDEr
N-Caltech101 [31] Qwen2-VL-Chat [38] 10B 0.346 0.153 0.069 0.033 0.096 0.287 0.092
DeepSeek-VL-Chat [30] 7B 0.217 0.096 0.042 0.022 0.113 0.233 0.034
LLaVA-v1.6 [29] 7B 0.294 0.148 0.073 0.039 0.105 0.269 0.098
Video-LLaVA [28] 7B 0.368 0.183 0.974 0.047 0.118 0.298 0.134
VideoLLaMA2 [9] 7B 0.393 0.200 0.107 0.061 0.121 0.326 0.150
EventVL-Base (Ours) 2.3B 0.360 0.238 0.163 0.115 0.145 0.377 0.823
HARDVS [40] Qwen2-VL-Chat [38] 10B 0.214 0.091 0.042 0.023 0.097 0.214 0.066
DeepSeek-VL-Chat [30] 7B 0.133 0.046 0.015 0.006 0.100 0.170 0.021
LLaVA-v1.6 [29] 7B 0.196 0.087 0.032 0.021 0.097 0.202 0.054
Video-LLaVA [28] 7B 0.299 0.144 0.087 0.092 0.103 0.257 0.124
VideoLLaMA2 [9] 7B 0.336 0.186 0.119 0.073 0.134 0.348 0.146
EventVL-Base (Ours) 2.3B 0.605 0.513 0.450 0.396 0.314 0.628 2.603
N-ImageNet [23] Qwen2-VL-Chat [38] 10B 0.197 0.087 0.040 0.021 0.067 0.208 0.152
DeepSeek-VL-Chat [30] 7B 0.161 0.071 0.033 0.017 0.092 0.191 0.100
LLaVA-v1.6 [29] 7B 0.240 0.119 0.061 0.034 0.083 0.203 0.186
Video-LLaVA [28] 7B 0.246 0.133 0.082 0.041 0.095 0.224 0.255
VideoLLaMA2 [9] 7B 0.319 0.167 0.094 0.057 0.106 0.273 0.358
EventVL-Base (Ours) 2.3B 0.561 0.392 0.279 0.205 0.167 0.462 0.573
Table 2: Quantitative performance of zero-shot event-based caption generation on different datasets. In this experiment, we only consider the event data as input for accurately evaluating the model’s event-based understanding ability. Our EventVL achieves competitive prediction performance compared with state-of-the-art MLLMs. Gray color denotes the video-based MLLMs.
Val set $n$ Bleu-4 METEOR ROUGE-L CIDEr
N-ImageNet 1 0.056 0.098 0.299 0.387
5 0.078 0.118 0.329 0.550
10 0.089 0.127 0.345 0.646
15 0.090 0.127 0.348 0.683
20 0.095 0.132 0.353 0.720
HARDVS 1 0.212 0.203 0.519 1.205
5 0.328 0.268 0.581 1.970
10 0.345 0.291 0.598 2.129
15 0.369 0.306 0.615 2.364
20 0.392 0.311 0.628 2.604
Table 3: Quantitative results about few-shot caption generation. “$n$” denotes the sample number per category used in training.
4.4 Inference
EventVL offers two kinds of inference pipelines. The first one is to only utilize event modality as input. The second type is to incorporate the image and event modality as input, in which we directly add event embeddings with image embeddings as the input embedding.
Figure 3: Qualitative results about zero-shot caption generation by utilizing N-ImageNet [23] and HARDVS [40] val set. Red color denotes model wrong description and orange color denotes the accurate object description. Underline text means redundant description.
5 Experiments
5.1 Experiments Setup
Implementation details can be referred to appendix.
Datasets. N-ImageNet [23] is the event camera version of ImageNet and the largest event camera dataset, which contains almost 1.2 million event streams and 1,000 classes. HARDVS [40] is a recently released dataset for event-based action recognition, currently having the largest action classes, namely 107,646 recordings for 300 action categories. DSEC [16] (Driving Stereo Event Camera) dataset is an autonomous driving dataset with a particular focus on event-based vision. We also utilize N-Caltech101 [31] for evaluating model zero-shot performance and further few-shot testing. Similar to N-ImageNet construction, N-Caltech101 contains 8,246 samples from 101 classes. More dataset details can be found in the appendix.
Dataset Preprocessing. Overall processed dataset information can be referred to Table 1. After we annotate these dataset by Section 3, we split the image/video-event-text pairs from N-ImageNet [23], HARDVS [40], N-Caltech101 [31] into the corresponding training sets and validation sets according to their categories. We use the whole DSEC training set by segmenting its scene video into multiple frame groups. The categories in each split do not overlap with each other in the same dataset for zero-shot evaluation. Few-shot datasets setting can be found in the appendix.
Figure 4: Qualitative results about zero-shot caption generation by utilizing DSEC [16] origin val set. Red color denotes model wrong description and orange color denotes the accurate object description.
5.2 Caption Generation
Zero-shot Evaluation. To evaluate the event-based understanding performance of our proposed framework, we conduct comparisons with the selected SOTA MLLM baselines, including image-based [38, 30, 29] and video-based MLLMs [28, 9]. As shown in Table 2, we can conclude that the video-based MLLMs can generate more high-quality descriptions for event data, which is determined by the stream properties of event data. Finally, experiment results demonstrate the effectiveness of EventVL to describe the event data. Our proposed EventVL significantly achieves 2.603 and 0.573 CIDEr in HARDVS and N-ImageNet respectively, outperforming existing works by a large margin, showing our advantages in event understanding ability. Furthermore, our model parameters are only 2.3B, proving its effectiveness, which is easy to deploy on the edge device.
Few-shot Evaluation. We also explore the impact of the training data scale on our event-based understanding performance. As shown in Table 3, the performance is closely related to the size of the dataset, which basically conforms to the principle of Scaling Law [10]. When $n=5$, the model learns more class invariant knowledge, which is beneficial for making more comprehensive perceptions. We also conclude that in a little data ($n\leq 10$), our model can get a significant promotion in event analysis, which verifies our design superiority. However, the improvement of event-based understanding ability is gradually slowing down when $n\geq 20$, because event-based knowledge learning has approached its limits in the limited parameters.
5.3 Qualitative results
We perform qualitative experiments on various scenario val sets with the zero-shot setting and utilize VideoLLaMA2 [9] to compare with our proposed EventVL. With only event data as input, as shown in Figure 3, our model can generate accurate descriptions for these event-based represenation, while VideoLLaMA2 always generate some useless text, like the denoted underline text “around in a field of dots”. The reason is that current MLLMs training data do not contain the event-based representation, which causes falling in explicit recognition wrong for factual hallucinations. In addition, our model can easily perceive objects and actions in hard scenes such as low light and blurry scenes, and provide comprehensive descriptions. Like the example in the upper left corner of HARDVS cases, our EventVL can not only describe this motion but also infer the colour of clothing worn by humans. Next, we also investigate the long description generation ability of EventVL by evaluating the drive scene. As shown in Figure 4, our proposed EventVL can generate a more precise description for the traffic scene and reasonable prediction ability like the second case “Traffic lights are green” in orange colour. However, VideoLLaMA2 easily has hallucinations when seeing these representations like the accident description in red color. Qualitative results show our advantage in understanding event-based presentation. Results about multiple rounds of dialogue and failure cases are also shown in the appendix.
Dataset Model Bleu-4 METEOR ROUGE-L CIDEr
N-ImageNet - 0.179 0.203 0.437 1.306
w/o ev. 0.218 0.233 0.471 1.572
w/o im. 0.052 0.098 0.252 0.203
N-Caltech101 - 0.313 0.255 0.562 0.874
w/o ev. 0.362 0.297 0.614 1.123
w/o im. 0.083 0.130 0.326 0.1110
HARDVS - 0.160 0.218 0.422 0.714
w/o ev. 0.227 0.242 0.499 1.219
w/o im. 0.012 0.099 0.189 0.040
Table 4: Exploration experiments of using the untraining EventVL initialized from InternVL2-2B [7]. “wo ev./im.” denotes that we don’t use the event or image modality as input.
Train Test Model Bleu-4 METEOR ROUGE-L CIDEr
+N-ImageNet N-Caltech101 - 0.435 0.289 0.636 1.273
w/o im. 0.180 0.162 0.444 0.377
[…]
w/o im. 0.031 0.126 0.243 0.089
[…]
w/o im. 0.121 0.151 0.386 0.889
+N-ImageNet +HARDVS N-Caltech101 - 0.425 0.287 0.634 1.275
w/o im. 0.176 0.161 0.445 0.372
[…]
w/o im. 0.386 0.317 0.618 2.522
[…]
w/o im. 0.117 0.148 0.379 0.833
+N-ImageNet +HARDVS +DSEC N-Caltech101 - 0.460 0.295 0.638 1.480
w/o im. 0.205 0.167 0.462 0.573
[…]
w/o im. 0.396 0.314 0.628 2.603
[…]
w/o im. 0.115 0.145 0.377 0.823
Table 5: Ablation study about the processed dataset choice for training by using event data as input. Note that categories from training sets aren’t overlapped with the test sets. “w/o im.” denotes that we don’t use the image modality as input.
5.4 Ablation Study
In this section, we utilize the zero-shot caption generation evaluation setting to perform the comprehensive ablation studies for investigating each effective module. More ablation studies can be found in the appendix.
Dataset Quality Assessment. In this part, we investigate the effectiveness of our proposed dataset. As shown in Table 4, we directly utilize the untraining EventVL initialized from InternVL2-2B to test on all validation sets for observing the current SOTA MLLM weights on the processed dataset performance. Experiments show that event representations transferred in the red-blue colour map make the MLLM confused, leading to serious understanding degradation. Furthermore, event data hinders the MLLM’s understanding of images. These observations prove event-based understanding deficiency of MLLMs and the urgency of our motivation. Meanwhile, we explore the enhancement of model training performance as shown in Table 5. We can get the conclusion that the single domain dataset training only makes a positive effect on the corresponding val sets while getting a tiny improvement on other val sets.
5.5 Extensive Application
Model Text2Event Event2Text
Recall@1 Recall@5 Recall@10 Recall@1 Recall@5 Recall@10
EventBind [52] 84.2 97.5 100 88.7 98.4 100
EventVL (Ours) 80.4 96.7 100 76.0 97.2 100
Table 6: Event retrieval results on processed N-Caltech101 [31].
We also explore other downstream applications based on our framework. For example, EventVL can be easily re-developed for event-image-text retrieval tasks. We utilize EventBind [52] for comparison as shown in Table 6. Experiment settings and discussion can be found in the appendix.
6 Conclusion
In this work, we introduce EventVL, the first event-based MLLM designed to achieve comprehensive understanding of event stream data. Through the integration of an advanced event encoder that captures rich semantic information, the novel Event Spatiotemporal Representation for efficient feature aggregation, and the Dynamic Semantic Alignment mechanism for fine-grained feature alignment, EventVL significantly advances the understanding and reasoning capabilities for event-based data. Additionally, we contribute a large-scale, high-quality annotated dataset with nearly 1.4 million event-image-text pairs, providing a robust foundation for future event-based research. We believe that EventVL sets a new direction for chat-centric event stream comprehension and lays the groundwork for future breakthroughs in event-driven multimodal systems.
Supplementary Material
7 Related Work
Event-based Vision. The microsecond-level temporal resolution, high dynamic range (typically 120 dB compared to the 60 dB of standard cameras), and power efficiency of event cameras represent a paradigm shift in imaging, moving beyond traditional frame-based methods [14]. Numerous event-based low-level imaging tasks, including recognition, perception, localization, and reconstruction, have been developed, covering areas such as object recognition [11, 33], object detection [44, 15], optical flow [12, 27], semantic segmentation [25], depth estimation [32], and object reconstruction [49, 22, 42, 13, 21], often utilizing techniques like NERF or Gaussian Splatting. While event-based models have demonstrated success in these traditional perception tasks, the exploration of event-based MLLMs (Multimodal Large Language Models) remains largely uncharted. In this paper, we focus on a novel application—event captioning and description for comprehensive event stream understanding. This endeavor aims to address the challenges posed by sparse, asynchronous, and high-temporal-resolution event data, with the goal of generating explicit language descriptions. Such capabilities are particularly crucial for safety-critical applications, such as drone or in-vehicle perception systems.
Figure 5: Problem list for different kinds of datasets. “Static Images” is for N-ImageNet [23] and N-Caltech101 [31], “Human Motions” is for [40] and “Drive Scenes” is for DSEC [16].
8 Preliminaries
The original stream output of an event camera consists of a set of asynchronously generated events, where each event is represented by $(x,y,t,p)$. Here, $(x,y)$ are the spatial coordinates, $t$ is the timestamp, and $p$ indicates the polarity of the intensity change: $+1$ for increased brightness and $-1$ for decreased brightness. These events feature a high dynamic range greater than 120 dB and a high temporal resolution equivalent to thousands of frames per second. However, directly processing event streams is challenging [14]. Previous works often represent events in frame-like representation, such as gray-style [43] or red-blue color map [25, 52], to simplify cross-modal alignment.
9 Data Engineering Setting
Algorithm 1 Generate Adaptive Ratios
Input: $n_{\text{min}},n_{\text{max}}$ are the min./max. number of patch.
Output: $\mathcal{K}$ is the adaptive ratio set.
1:$\mathcal{K}\leftarrow\{\}$
2:for $n\leftarrow n_{\text{min}}$ to $n_{\text{max}}$ do
3:    for $i\leftarrow 1$ to $n$ do
4:       for $j\leftarrow 1$ to $n$ do
5:          if $i\times j\leq n_{\text{max}}$ and $i\times j\geq n_{\text{min}}$ do
6:             $\mathcal{K}\leftarrow\mathcal{K}\cup\{(i,j)\}$
[…]
8:return $\mathcal{K}$
In this section, we mainly introduce the origin dataset information, the problem list and process details.
9.1 Origin Dataset
N-ImageNet [23]. As the largest event camera dataset, it contains 1.78 million event streams and 1,000 classes. The data were captured with a moving 480 × 640 resolution Samsung DVS Gen3 event camera under the same environmental conditions. N-ImageNet also provides variants of the test set captured with different camera motions and brightness, serving as a benchmark to evaluate the robustness of event classifiers. The authors create 9 variants of the same test set. Variants 1-5 change the camera motions used to trigger events, where different moving directions, frequencies, and amplitudes of the camera trajectory are employed. Variants 6-9 alternate the lighting conditions of the environment, such as extremely low or high illuminations. Overall, these variations cause a large degradation in the performance of existing event-based classifiers.
N-Caltech101 [31]. N-Caltech101 contains 8,246 samples from 101 classes, recorded by a moving 180 × 240 resolution ATIS system in front of a monitor displaying still images from the original RGB Caltech101 dataset.
HARDVS [40]. HARDVS is the first large-scale realistic neuromorphic human activity recognition dataset. Specifically, the HARDVS dataset contains more than 100K video clips recorded with a DAVIS346 camera, each of them lasting for about 5-10 seconds. It contains 300 categories of human activities in daily life, such as drinking, riding a bike, sitting down, and washing hands. The following factors are taken into account to make HARDVS data more diverse, including multi-views, illuminations, motion speed, dynamic background, occlusion, flashing light, and photographic distance.
DSEC  [^1] https://dsec.ifi.uzh.ch/ [16]. DSEC is an extensive dataset designed for advanced driver-assistance systems (ADAS) and autonomous driving research, with a particular focus on event-based vision and stereo vision. Different from DDD17 [3], DSEC dataset combines data from event-based cameras and traditional RGB cameras. Including event-based cameras (which capture changes in light intensity) alongside regular cameras provides a rich, complementary data source for perception tasks. The dataset typically features high-resolution images and event data, providing detailed visual information from various driving conditions, including urban, suburban, and highway environments, various weather conditions, and different times of the day. This diversity is crucial for developing systems that can operate reliably in real-world conditions. In this paper, we only utilize the training set of the DSEC for post-processing. In detail, we first split each scene video into multiple segments by 24 frames per segments. Then we utilize these segment for generating our description by Section 3.
9.2 Processing details
Our different problem lists are as shown in Figure 5. In each round annotating the data, we randomly sample the problem from the corresponding lists. For N-ImageNet, we directly utilize the whole event-pairs dataset for training, and we do not use its variants. For N-Caltech101, we utilize it only to validate the effectiveness of our proposed EventVL. For HARDVS, we directly utilize its red-blue color map representation instead of the raw event data for efficiency. In each category, we uniformly sample 14 frames for each video for post-processing. If videos contain less than 14 frames, we directly utilize the whole video. For DSEC, we only sample 14 frames in each segment for training efficiency.
Figure 6: Case visualization of the proposed Event Spatiotemporal Representation.
10 Methods
As shown in Figure 6, we give a case to visualize the event stream processed by our proposed Event Spatiotemporal Representation. We decouple the event stream from time and space to obtain a comprehensive representation of the event stream. Then, We give more explanation of the proposed Adaptive Spatial Split. For predefined set $\mathcal{K}$, we utilize Algorithm 1 to complete the $\mathcal{K}$. As we set $n_{min}$ to 1 and set $n_{max}$ to 6, we have $\mathcal{K}\in${1:1,1:2,…,6:1}. Then, we utilize the $\mathcal{K}$ for following event spatial segmentation in Section 4.1.
11 Experiments
11.1 Implementation
Fundamental Module. We selected the InternViT-300M-448px  [^2] https://huggingface.co/OpenGVLab/InternViT-300M-448px [7] as the foundational architecture for both the image and event encoders, given its proven effectiveness in various image and video understanding benchmarks. This model contains 24 attention blocks, as shown in Figure 7. The term “Drop” refers to the drop path, commonly applied in the main path of residual blocks. In our setup, $B$ denotes the batch size, $C$ represents the channel number, and $H$ and $W$ are the height and width of the image or event. $P$ is the patch number, and $C_{\phi}$ is the final output embedding channel number. We set $H=W=448$, $C_{\phi}=2048$, and $P=256$. We also leverage the pretrained weights of InternViT for initialization, allowing us to transfer spatial prior knowledge from images to events, which accelerates training and improves data efficiency for EventVL. For the text decoder, we chose InternLM2-1.8B [4], which has shown significant improvements in capabilities such as reasoning and coding.
Figure 7: Illustration of the encoder architecture including image or event encoder. The encoder contains 24 attention blocks.
Figure 8: Qualitative results about processed zero-shot caption generation by utilizing N-ImageNet [23] val set. Red color denotes model wrong description and orange color denotes the accurate object description. Underline text means redundant description.
General Setting. In Dynamic Semantic Alignment, we set $n_{min}=1$ and $n_{max}=6$. We also set the $\lambda_{1}=1$ and $\lambda_{2}=1$ for balancing the Multimodal alignment. Caused by the difference between the event number scale between N-ImageNet and N-Caltech101, we set different event point numbers per frame $n_{\epsilon}$ for each dataset, which we set $n_{\epsilon}=40000$ for N-ImageNet and $n_{\epsilon}=20000$ for N-Caltech101. If the total event point number $N_{\epsilon}$ does not fulfill the demand of all level event-based representation from Section 4.1, we pad this event stream for post-processing. Note that we do not utilize Event Spatiotemporal Representation for video-event pairs from HARDVS [40] and DSEC [16] datasets, which contain enough temporal and spatial information. We directly utilize cosine similarity from Equation 1 to align the video-event modalities. The final optimization objectives still are the same as the image-event pairs datasets training. Due to the unbalanced data scales between these domains, we adopt the weighted sampling strategy for hybrid Training. We set the training sampling weight of 0.6 for N-ImageNet, 0.1 for DSEC, and 0.3 for HARDVS. COCO caption metrics [6] is used for quantitative analysis model performance. All training is performed on eight Nvidia GTX A6000 GPUs for approximately 120 hours. We set the batch size to 2 and adopt the DeepSpeed  [^3] https://github.com/microsoft/DeepSpeed ZeRO-1 training strategy for training efficiency.
Zero-shot Setting. We utilized N-ImageNet, HARDVS and DSEC train splits for training. To overcome overfitting, we adopt 1 epoch for training. We evaluate EventVL on the val sets for zero-shot description generation with only event data as input.
Few-shot Setting. For few-shot experiments, we sample $n$-shot data from the all data of N-ImageNet, HARDVS, and DSEC for few-shot training. Then, we utilize the remaining datasets including N-Caltech101 to evaluate its performance. We adopt 3 epochs to perform all few-shot training in order to learn enough event-based category knowledge.
L1 L2 L3 N-Caltech101 HARDVS N-ImageNet
ROUGE-L CIDEr ROUGE-L CIDEr ROUGE-L CIDEr
- - 0.325 0.769 0.582 1.885 0.434 0.548
- 0.349 0.789 0.598 2.409 0.445 0.554
- 0.353 0.786 0.601 2.411 0.441 0.557
0.377 0.823 0.628 2.603 0.462 0.573
Table 7: Different level of the event-data aggregation choice for event batch construction. “L1/L2/L3” denotes the Level-1/2/3.
11.2 More Cases For Zero-shot Evaluation
Figure 9: Qualitative results about processed zero-shot caption generation by utilizing HARDVS [40] val set. Red color denotes model wrong description and orange color denotes the accurate object description.
As shown in Figure 8 and Figure 9, we provide additional examples to validate the superiority of our proposed EventVL framework using only event data as input. These examples demonstrate that EventVL captures more fine-grained details from the given event stream compared to VideoLLaMA2, highlighting the effectiveness of our framework.
Figure 10: Qualitative results about multi-round dialogue by utilizing N-ImageNet [23] val set. Blue color denotes the detailed description.
11.3 Multi-round Dialogue
As illustrated in Figure 10, we present several examples demonstrating the application of our proposed EventVL framework in multi-round dialogue based on event data. These examples reveal that EventVL can understand comprehensive semantics within event data while also capturing finer object details, such as describing “a clear blue sky”.
$n_{max}$ N-Caltech101 HARDVS N-ImageNet
ROUGE-L CIDEr ROUGE-L CIDEr ROUGE-L CIDEr
1 0.310 0.756 0.565 2.398 0.431 0.528
5 0.341 0.810 0.603 2.575 0.459 0.569
6 0.377 0.823 0.628 2.603 0.462 0.573
7 0.388 0.841 0.633 2.648 0.473 0.592
8 0.386 0.843 0.635 2.649 0.468 0.575
Table 8: Different $n_{max}$ settings for Dynamic Semantic Alignment. Note that we default the $n_{min}=1$.
11.4 Ablation Study
Dataset Quality Assessment. We investigate the impact of training on individual datasets to enhance model performance, as summarized in Table 5. Our findings indicate that training across multiple domains improves the model’s general event-based understanding by mitigating overfitting issues seen in single-domain training. For example, this approach yields nearly a 20% increase in CIDEr on the N-Caltech101 dataset. Additionally, incorporating scene data (DSEC) enhances the model’s ability to capture broader contextual details, evidenced by CIDEr gains on HARDVS compared to training without DSEC. Moreover, improvements in event-based understanding bolster the model’s overall multimodal comprehension. Compared to results in Table 4, our method achieves 1.769 CIDEr (+46.3%) on the N-ImageNet validation set, significantly outperforming the untrained model (1.306 CIDEr).
Figure 11: Illustration of the failure cases of our proposed EventVL. Red color denotes the wrong description.
Event Representation. As shown in Table 8, we observe that excluding the dynamic patch setting ($n_{max}=1$) results in significant performance degradation. Lower values of $n_{max}$, such as $n_{max}=5$, lead to coarse alignment between images and events due to insufficient semantic information. Conversely, increasing $n_{max}$ to values like $n_{max}=8$ does not enhance EventVL’s understanding, as its category knowledge reaches inherent limits. Additionally, larger values of $n_{max}$ consume more CUDA memory. To balance performance and resource consumption, we select $n_{max}=6$ for efficient temporal interaction and semantic extraction. We also examine various levels of aggregation in the Adaptive Spatial Split process. As shown in Table 7, results indicate that all-level event aggregation provides the best performance, while any omission of representation at any level causes a decline in our framework’s effectiveness.
Optimization Setting. As shown in Table 9, we first analyze the main components of our training objectives. When using only cross-entropy for event-text alignment, the model performance decreases by 9.7% in CIDEr on the N-ImageNet evaluation. Incorporating the image prior improves performance, likely by bridging semantic gaps between text and event data. The proposed $\mathcal{L}_{c}$ introduces an additional 5.8% improvement in CIDEr on N-ImageNet, enhancing multimodal alignment.
We also examine the impact of varying weight settings in the optimization objectives of Equation 5. As shown in Table 10, the results indicate that text-based alignment exerts a greater influence on the model’s understanding of event data. Lower values of $\lambda_{1}$ correlate with a more significant drop in performance, demonstrating its importance in our optimization strategy.
Prior $\mathcal{L}_{c}$ N-Caltech101 HARDVS N-ImageNet
ROUGE-L CIDEr ROUGE-L CIDEr ROUGE-L CIDEr
- - 0.315 0.725 0.524 1.996 0.324 0.466
- 0.342 0.747 0.568 2.394 0.358 0.491
- 0.356 0.734 0.545 2.373 0.334 0.524
0.377 0.823 0.628 2.603 0.462 0.573
Table 9: Various loss construction for training objective Equation 5. “Prior” denotes the image prior and $\mathcal{L}_{c}$ denotes the image-event embedding similarity loss.
11.5 Failure Cases
$\lambda_{1}$ $\lambda_{2}$ N-Caltech101 HARDVS N-ImageNet
ROUGE-L CIDEr ROUGE-L CIDEr ROUGE-L CIDEr
0.2 1 0.298 0.368 0.512 1.235 0.324 0.398
0.5 1 0.319 0.440 0.553 1.549 0.330 0.435
1 0.1 0.291 0.667 0.467 1.874 0.379 0.317
1 0.2 0.298 0.689 0.488 2.102 0.387 0.341
1 0.5 0.345 0.724 0.574 2.263 0.413 0.451
1 1 0.377 0.823 0.628 2.603 0.462 0.573
Table 10: Different $\lambda_{1}$ and $\lambda_{2}$ setting for the training objective of Equation 5.
As shown in Figure 11, we present several failure cases of our proposed EventVL. We summarize three types of errors that occur in EventVL as follows: Misclassification (Case (a,f,g)), Misinterpretation (Case (d,e,i)) and Recognition Failures (Case (b,c,h)). For Misclassification, our model limits in distinguishing between different animal categories when facing the poor event stream quality. For instance, in case (a), the model misidentifies a chicken as “a dog”. Second, for Misinterpretation, our model fails to accurately recognize ongoing activities or scene elements such as case (d), where an elderly person playing a harp is mistakenly described as “painting a window”. Even though multimodal alignment forces the model to learn such knowledge in post-processing, the model still falls short in capturing event data surface details such as object textures. Finally, we also observe the Recognition Failures. For example, in case (c), the model fails to mention the prominent pink flowers in the scene. These failure cases reveal a fundamental challenge in applying vision-based models to event camera data. The distinctive texture patterns and unique visual representations generated by event cameras can introduce ambiguity that conventional vision models struggle to interpret accurately. In the future, we will explore more efficient strategy to bridge this gap between images and events for better event-based understanding.
11.6 Event Retrieval
The complex architecture of decoder-only LLMs makes it challenging to adapt them as text encoders. Recent research has attempted to modify the inherent mechanisms of LLMs to handle text representation tasks [2], but this approach often lacks efficiency. Inspired by [20], we instead freeze the overall model and add trainable attention layers to both the encoder and decoder, followed by contrastive learning. This approach allows for efficient retrieval tasks with strong performance. Specifically, we first add an attention layer to each frozen encoder, and a linear projection layer to the text decoder. These additional layers are trainable, designed to reduce the high-dimensional complexity of the model to a lower-dimensional space for more efficient alignment. We then apply contrastive learning [5] to achieve significant semantic alignment across modalities. We set the training epoch to 3, keeping other parameters consistent with previous experiments. As shown in Table 6, our proposed strategy achieves comparable performance to EventBind, demonstrating the extensibility of our EventVL framework.
12 Discussion
Training Methods. In this section, we discuss the various training approaches employed in our methods, including full fine-tuning, LoRA [19], and LLM Adaptors [20]. LoRA is commonly used in contexts with limited computational resources and small datasets. We initially experimented with LoRA for training our framework on a small batch of data, but found that its multiple manually-tuned parameters increased the risk of training instability within our framework. We also opted not to use adaptor-based training, as adaptors consist of learnable parameters tied to the original modality, which can limit model performance. Event data, containing more comprehensive information about the captured scenes than traditional camera data [37], requires a large parameter space to adequately capture this information. Given our access to abundant, high-quality paired data, we selected full fine-tuning as the primary training method for our framework. This approach effectively captures event-based categorical knowledge, supporting a deeper understanding of event-based world semantics.
Scale. Our processed datasets span a variety of domains, including static images [23, 31], body motion [40], and driving scenes [16]. However, the data remains insufficient in certain areas, such as sports movements and complex scenes, which limits the model’s depth of understanding of event data. This is particularly evident in the driving domain, where high-resolution event data is still lacking, hindering the clarity of perception and collaborative development between event and regular cameras. To address this, we aim to construct a larger and more comprehensive dataset. Additionally, due to computational resource limitations, we are constrained to using models with fewer than 3 billion parameters for training or fine-tuning. We believe that our proposed framework can be readily scaled to models with a larger parameter count, enhancing its capacity for event-based understanding.
Event Representation. In line with the current VLM setting [46, 48, 25], we use a red-blue color map event representation for processing and training. While this approach is effective, it may not be optimal for an MLLM-based framework. In future work, we plan to further investigate the impact of different event representations [41, 50] to explore alternatives that may better support our framework’s performance.